# PinRec: Outcome-Conditioned, Multi-Token Generative Retrieval for Industry-Scale Recommendation Systems

### v4

@@ 4.3. Training Objective
During training, we employ teacher forcing with a next-item prediction objective. The hidden states $\mathbf{h}_{u,t}$ are processed through the conditioned output head $O$ to generate item representations that predict future items in the sequence. For outcome-conditioned generation, we condition the output head on the actual values of the target item during training. For multi-token prediction, we sample $k$ positions from the next $N$ positions in the sequence to serve as prediction targets.
Following Pinnerformer22, we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (cormode2005improved) to account for the non-uniform sampling of in-batch negatives.
@@
We average this loss over all positions in each user sequence. For standard next-token prediction, the target is simply the immediate next item:
(4) $L_{\theta}^{\text{next}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
However, this strict sequential assumption may not align well with user interaction patterns. User interaction sequences differ fundamentally from language: while they exhibit global structure reflecting long-term interests, they lack the local structural constraints found in natural language. On social media platforms, users engage with feed suggestions in arbitrary order within sessions, making strict positional dependencies less meaningful than in natural language modeling.
To better accommodate these interaction patterns, we introduce an intra-feed relaxation that considers all items engaged within the same feed session as valid targets. Given a feed session $F_{s}=\{i_{t_{1}},i_{t_{2}},\ldots,i_{t_{n}}\}$ where $t_{1}<t_{2}<\ldots<t_{n}$, we define the relaxed target set for position $t_{j}$ as:
$T_{t_{j}}^{\text{feed}}=\{i_{t_{k}}\mid t_{k}\in F_{s},t_{k}>t_{j}\}$
This enables a relaxed objective that optimizes for predicting any future item within the same session:
(5) $L_{\theta}^{\text{feed}}(H(u,t_{\mathrm{max}}))=\frac{1}{|H(u,t_{\mathrm{max}})|}\sum_{t_{j}}\min_{i\in T_{t_{j}}^{\text{feed}}}L_{s}(\hat{\mathbf{i}}_{u,t_{j}},i)$
where we minimize the loss to the best-matching item among all future items within the same feed session, rather than strictly requiring prediction of the immediate next item. Intuitively, this optimization can be posed as: can we predict any of the future items that the user will engage with within the same feed session? We find that this approach is more naturally aligned with user interaction patterns on Pinterest compared to strict next-token prediction.
Finally, we minimize the expected loss across all user sequences in dataset $\mathcal{D}$:
(6) $\arg\min_{\theta}\mathbb{E}_{H(u,t_{\max})\sim\mathcal{D}}[L_{\theta}^{\text{feed}}(H(u,t_{\max}))].$
@@ 4.4. Inference
To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration:
Unconditional Generation. For the baseline model, generation is straightforward: at each step, the model produces a single item embedding, which is appended to the sequence along with the target surface embedding and a predefined temporal offset (e.g., 10 seconds). This updated sequence is then used to generate the next embedding.
Outcome-Conditioned Generation. For models with outcome conditioning, for each outcome we generate a candidate embedding, producing multiple candidate embeddings per generation step. For autoregressive continuation, we sample a value for each condition according to the requested distribution over each condition $D_{1},D_{2},\ldots,D_{n}$ and use the corresponding embedding as input for the next step (Fig. 2).
Temporal Multi-Token Generation. When using temporal multi-token prediction, we sample time deltas up to a specified maximum and generate multiple embeddings at different temporal offsets, combined with any specified outcome conditions. For the next autoregressive input, we use the embedding corresponding to the immediate next token.
Across $n$ rollouts, this process generates varying numbers of embeddings depending on the model configuration and input settings. The generated embeddings are then used for candidate retrieval through Approximate Nearest Neighbor (ANN) search.
@@
This candidate generation process enables PinRec to serve as a flexible first-stage retrieval system within Pinterest’s recommendation infrastructure. The generated candidates can be further refined by downstream ranking models, allowing the system to benefit from both the generative model’s ability to capture complex user patterns and the precision of specialized ranking algorithms. The outcome-conditioning and multi-token capabilities make the system particularly suitable for multi-objective optimization scenarios common in production recommendation systems.
@@ 4.5. Serving
PinRec is implemented with an in-house, optimized implementation of transformer and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval. This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with learned ID embeddings, autoregressive inference for item generation, and Faiss retrieval (Figure 3).
@@ 4.5.1. Signal Fetching
Our Signal Service uses a Lambda architecture that combines both batch and real-time processing. The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, storing the sequences and features for each user ID. The real-time component is backed by a RocksDB key-value store (kvstore) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time. The batch and real-time signals are deduplicated and returned to PinRec. On surfaces with context (e.g., Search, Related Pins), model inference also includes the context (e.g., search query, closeup Pin) issued by the user at request time as the final element in the sequence. To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision and later dequantized to FP16 for inference.
4.5.2. Learned ID Embeddings
As discussed in Section 4.2, our pin embedder includes an ID embedding table that maps each pin ID’s hash to an embedding, followed by MLP layers. Since this ID embedder requires significant memory, especially when replicated, it runs as a separate CPU memory-optimized service to avoid competing with the main model for GPU resources. It is served with torchscript to achieve single-digit millisecond latencies.
@@ 4.5.4. Faiss Retrieval
(a) Recent user history specified to produce results in Subfigure 4(b).
(b) Results for the user sequence specified in Subfigure 4(a).
Figure 4. Item recommendations from PinRec-OC for a user history with cat-related search queries and pin interactions (history shown in reverse chronological order). Additional examples are provided in Section A.2.
@@
Sample results from this complete serving pipeline are visualized for a manually specified user history in Figure 4.
@@ 5. Offline Evaluation
In this section, we dissect the performance of the proposed generative recommendation system, PinRec, on generative retrieval across a range of offline evaluations. For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions) with an 80M-row ID embedding table (256 total dimensions, 8 dimensions per hash). Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Homefeed, Related Pins, and Search. We demonstrate that outcome-conditioned generation allows for improved controllability and that multi-token generation leads to significantly improved diversity, performance, and efficiency.
@@ 5.1. Offline Evaluation Metrics
We construct the evaluation set using interactions from users disjoint from the training set and evaluate whether the proposed technique can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”). For such retrieval tasks, recall@$k$ is a standard performance metric, measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
We evaluate model performance using a sequence generalization of recall@$k$ that accounts for the multi-embedding nature of our autoregressive generation. Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of predictions to sets of targets for each user.
@@
We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on any target items during prediction, making the evaluation more realistic for deployment scenarios.
Beyond retrieval performance, we examine recommendation diversity by computing the proportion of unique items retrieved across all users in the evaluation set after selecting top-$k$ predictions. This quantifies whether learned representations exhibit bias toward popular items versus providing diverse recommendations across the user base.
5.2. Offline Results
In this section, we compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pinnerformer22), SASRec (kang2018self), and the leading generative recommendation method TIGER (Tiger23). We also evaluate our transformer architecture against the proposed HTSU (HSTU24) module. Furthermore, we quantify the impacts of our design choices on both recommendation performance and diversity.
5.2.1. Comparison with Prior Method
We evaluate three different variants of PinRec.
  • PinRec-UC: We do not condition on anything (i.e., $\Phi=\emptyset$) and do not employ multi-token prediction.
  • PinRec-OC: We condition on the target action (“outcome conditioned”) and do not employ multi-token prediction.
  • PinRec-{MT, OC}: We condition on the target action and employ multi-token prediction.
To ensure that the comparison between PinRec-{MT, OC} and other methods is fair, we adjust the number of autoregressive generation steps for multi-token generation to ensure that the total number of generated embeddings is held constant. As in training (where we cannot modify the evaluation dataset post-hoc), we condition on the target item’s action for outcome-conditioned variants. For simplicity, we reproduce results for the outcome-conditioned variant only when comparing against HTSU (additional details in Section A.1.1). The impact of negative sampling and model scaling effects are analyzed in Section A.1.2 and Section A.1.3, respectively.
Our comparison in Table 1 shows that PinRec outperforms other baselines across the board. Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Homefeed and Related Pins). Above all, PinRec-{MT, OC} shows improvements over every method, with over +10% lift on Homefeed, over +20% on Related Pins, and +30% on search over PinRec-UC.
@@ 5.2.1. Comparison with Prior Method
SASRec (kang2018self) 0.382 0.426 0.142
PinnerFormer (Pinnerformer22) 0.461 0.412 0.257
TIGER (Tiger23) 0.208 0.230 0.090
HTSU-OC (HSTU24) 0.596 0.539 0.179
@@
PinRec-{MT, OC} 0.676 0.631 0.450
5.2.2. Outcome-Conditioned Generation
@@
We assess whether conditioning on a desired action (e.g., “repin”) leads to outputs that elicit that action. If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action and lower recall for others. As a baseline, we compare against the unconditioned PinRec variant.
For simplicity, we allocate the entire budget to one action and perform offline evaluation conditioning across each action. We show the result of this analysis in Figure 5. Whenever we condition on an action outcome, we show notable gains over an unconditioned baseline for that action. Specifically, matching outcome conditioning yields 1.9% recall lift for true repins, 1.4% on true grid clicks, 4.8% on long grid clicks, and 6.2% on outbound clicks.
On the other hand, we observe decreases in recall compared to an unconditioned baseline if the true action does not correspond to the desired action, indicating that our conditioning achieves the desired result of producing outcome-specific representations. Especially for outbound clicks, we observe that conditioning on any other actions leads to massive drops in unordered recall.
Figure 5. Percentage lift in unordered recall for PinRec-OC over PinRec-UC when conditioning on the desired action, stratified by the actual action taken by the user.
5.2.3. Utility of Multi-Token Generation
We evaluate the benefits of multi-token generation as a function of the number of generations produced per timestep. Specifically, we analyze the tradeoff between generating more representations in fewer steps versus fewer representations over more steps, while keeping the total number of output embeddings constant. Since the strongest offline performance aside from the multi-token variant is by PinRec-OC, we consider it as a baseline. For simplicity, we average the unordered recall @ 10 across the three surfaces.
We show the tradeoff between performance, diversity, and efficiency with the number of embeddings we generate in each generation step using multi-token generation in Figure 6. While next-token generation (i.e., 1 embedding per step) is the most inefficient and yields the poorest performance (where we perform similarly to PinRec-OC), we observe significant benefits across the board in increasing the number of embeddings per step. At 16 generations per step, we demonstrate a $\sim$10x reduction in latency over PinRec-OC, with improved unordered recall (+16.0%) and diversity (+21.3%).
Figure 6. Tradeoff between unordered recall @ 10 (top), proportion unique (bottom), and generation latency (right) with the number of multi-token generations per step.
@@ 6. Online A/B Experiments
We perform online A/B experiments to evaluate PinRec as a candidate generator, measuring its impact on user engagement and confirming its online controllability. The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other candidate generators, to produce the final results shown to users. We tested on the three major surfaces: Homefeed, Search and Related Pins.
6.1. Homefeed
6.1.1. Experiment Configuration
In the Homefeed online experiment, we run five groups exposed to global traffic for PinRec-OC, with varying action budgets $\mathcal{B}$. The budgets are fixed for outbound click (0.1), long grid clicks (0.05), and other unspecified actions (0.05). The remaining budget is split between the repin and grid click engagements, in steps of 0.2 ranging from 0 to 0.8. We similarly configure five groups for PinRec-{OC, MT}, with the same budgets. We include control groups without PinRec.

### v5

@@ 4.3. Training Objective
As described in Sec. 1, PinRec is trained in two stages: pretraining and fine-tuning. Below, we formulate the training objective for each stage. Both stages use the same softmax-based loss; they differ only in the target outputs used for supervision and negative examples.
During both training stage (pretraining and fine-tuning), we employ teacher forcing with a next-item prediction objective. The hidden states $\mathbf{h}_{u,t}$ are processed through the conditioned output head $O$ to generate item representations that predict future items in the sequence. For outcome-conditioned generation, we condition the output head on the actual values of the target item during training.
Following Pancha et al. (2022), we compute similarity between the predicted item representation $\hat{\mathbf{i}}_{u,t}$ and candidate item $\mathbf{i}_{c}$ as $s(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{c})=\lambda\cdot\hat{\mathbf{i}}_{u,t}^{\top}\mathbf{i}_{c}-Q(\mathbf{i}_{c})$, where $Q(\mathbf{i}_{c})$ is a bias correction term estimated using count-min sketches (Cormode and Muthukrishnan, 2005) to account for the non-uniform sampling of in-batch negatives.
@@
Pretraining. We pretrain the model with next-token prediction on the vast amount of life long user history sequence $H(u,t_{\text{max}})$ (Sec. 3), which contains actions from across all surfaces in Pinterest and excludes impression-only items (Fig. 2). We compute the loss in Eq. 2 at each position in the user history sequence and average it across positions.
(3) $L_{\theta}^{\text{pre}}(H(u,t_{\max}))=\frac{1}{|H(u,t_{\max})|}\sum_{t=1}^{t_{\max}-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
Fine-tuning. For surface-specific fine-tuning, we collect the feed session sequence $F(u,s,t)$ for user $u$ on surface $s$ at time $t$, together with the user’s history up to time $t$ to avoid leakage, $H(u,t)$ (Sec. 3). To perform next-action prediction over both contexts, we construct a combined activity sequence $G$ by concatenating $H(u,t)$ with the action-only entries from $F(u,s,t)$:
[…]
We then compute the sequence loss on $G(u,s,t)$ in the same way as in pretraining, averaging the per-position losses.
(4) $L_{\theta}^{\text{ft}}(G(u,s,t))=\frac{1}{|G(u,s,t)|}\sum_{t=1}^{|G|-1}L_{s}(\hat{\mathbf{i}}_{u,t},\mathbf{i}_{u,t+1})$
Additionally we add impression-only items in $F(u,s,t)$ to the in-batch negative set $\mathbf{N}$ to serve as hard negatives (Fig. 2). We show in Sec. 5 that these impression-based negatives improve performance.
@@ 4.4. Inference
To generate an ordered set of item recommendations, we employ autoregressive generation on the transformer. The specific generation process varies based on the model configuration (Fig 3):
Unconditional Generation. Without outcome conditioning, generation is straightforward: at each step, the model produces a single item embedding, appended to the sequence along with the target surface embedding and a predefined temporal offset $o=10s$. This updated sequence is then used to generate the next embedding.
Outcome-Conditioned Generation. We generate one candidate embedding per outcome at each step, yielding multiple candidates per generation step. The overall FLOPs increase only marginally because a single forward pass can produce embeddings for multiple outcomes, similar to a multitask model. For autoregressive continuation, we sample one outcome from the set of considered outcomes and feed its corresponding embedding into the next step (Fig. 3).
@@
This candidate generation process enables PinRec to serve as a flexible first-stage retrieval system within Pinterest’s recommendation infrastructure. The generated candidates can be further refined by downstream ranking models, allowing the system to benefit from both the generative model’s ability to capture complex user patterns and the precision of specialized ranking algorithms. The outcome-conditioning makes the system well suited to the multi-objective optimization needed to customize recommendations for each surface.
@@ 4.5. Serving
PinRec uses an in-house optimized transformer implementation and served with NVIDIA Triton Inference Server’s ensemble setting and Python backend to perform autoregressive generative retrieval. This multi-stage serving pipeline includes separate stages for fetching signals, featurizing Pins with Pin features, autoregressive inference for item generation, and Faiss retrieval (Figure 4).
@@ 4.5.1. Signal Fetching
Our Signal Service uses a Lambda architecture that combines both batch and real-time processing. The batch signal is updated daily through a Spark pipeline that processes the past year’s positive user engagements and search queries, storing the sequences and features for each user ID. The real-time component is backed by a RocksDB key-value store (Prasad, 2024) that efficiently retrieves user engagements occurring after the batch signal’s cutoff time. The batch and real-time signals are deduplicated and returned to PinRec. On surfaces with context (e.g., Search, Related Pins), model inference also includes the context (e.g., search query, closeup Pin) issued by the user at request time as the final element in the sequence. To limit transport costs, all embeddings associated with sequence elements are quantized to INT8 precision and later dequantized to FP16 for inference.
@@ 4.5.3. Faiss Retrieval
4.5.4. Serving Efficiency
On the memory front, PinRec uses about 1.6GiB of GPU memory including KVCache and CUDA graphs, without any significant memory usage increase for outcome conditioning. Excluding the KVCache and CUDA graphs, we use roughly the same amount of memory ($\approx$250 MiB) relative to existing two-tower retrieval.
In terms of latency, at 80 QPS, PinRec-OC’s latency was p50 40ms/p90 65ms, and PinRec-UC’s was p50 23ms and p90 46ms. Though the latency is roughly 3-4x higher (p50) than a traditional two-tower solution, we note that parallelizing it with other calls results in ¡1% added end-to-end increase in latency.
@@ 5. Offline Evaluation
In this section, we dissect the performance of PinRec on generative retrieval across a range of offline evaluations. For the results presented, PinRec uses a 12-layer transformer model (12 attention heads, 768 hidden dimensions). Our experimental setting includes interactions on three major surfaces for the Pinterest platform, namely Home Feed, Related Pins, and Search. We demonstrate that outcome-conditioned generation captures user intent more accurately, and that surface-specific fine-tuning further improves metrics on each surface. Importantly, these two techniques are complementary: neither makes the other redundant. Each provides additional gains, and their benefits accumulate when used together.
@@ 5.1. Offline Evaluation Metrics
We construct the evaluation set using user activities disjoint from the training set and evaluate whether models can produce representations that retrieve future item embeddings from a set of $1M$ random items (the ”index”). We use recall@$k$ (Pancha et al., 2022; Ying et al., 2018; Yang et al., 2024), a standard performance metric measuring the proportion of target items correctly retrieved within the top-$k$ ranked items.
We evaluate model performance using a sequence generalization of recall@$k$ that accounts for the multi-embedding nature of our autoregressive generation. Unlike standard recall, which compares individual predictions to individual targets, our unordered recall metric compares sets of autoregressive generations to sets of future targets for each user.
@@
We perform autoregressive generation without teacher forcing, ensuring the transformer is not conditioned on target items during prediction, making evaluation realistic for deployment scenarios.
@@
SASRec (Kang and McAuley, 2018b) 0.382 0.426 0.142
PinnerFormer (Pancha et al., 2022) 0.461 0.412 0.257
TIGER (Rajput et al., 2023) 0.208 0.230 0.090
HSTU (Zhai et al., 2024) 0.596 0.539 0.179
@@
5.2. Offline Results
In this section, we compare PinRec against existing models and quantify the impacts of outcome-conditioning, pretraining and fine-tuning.
5.2.1. Model Evaluation
We compare PinRec against existing sequence-based user modeling approaches, including Pinnerformer (Pancha et al., 2022), SASRec (Kang and McAuley, 2018b), and the leading generative recommendation method TIGER (Rajput et al., 2023) and HSTU (Zhai et al., 2024). We evaluate different variants of PinRec.
  • PinRec-UC: We do not condition on anything (i.e., $\Phi=\emptyset$)
  • PinRec-OC: We condition on the target action (“outcome conditioned”)
To make it easier and fair to evaluate many model architectures, we use the same training procedure for all methods; all models are trained from the activity sequence from all surfaces like PinRec pretraining (Eq. 3). We evaluate on user activity sequences from held-out users (i.e., users not seen during training). There is no fine-tuning process involved in this evaluation. Additional details for HSTU evaluation, the impact of negative sampling and model scaling effects are in Sec. A.1.1, A.1.2 and A.1.3, respectively.
Our comparison in Table 1 shows that PinRec outperforms other baselines across the board. Importantly, outcome-conditioning demonstrates notable improvements over PinRec-UC (unconditioned) in unordered recall @ 10 (e.g., +2-3% on Home Feed and Related Pins).
5.2.2. Fine-tuning Evaluation
We evaluate the impact of fine-tuning using surface-view logs. Since surface traffic and content evolve over time, we adopt a time split: we train on logs before a cutoff time and evaluate on logs collected after the cutoff, following standard industrial recommender evaluation protocols (Chen et al., 2025; Xia et al., 2023). This time-split evaluation set differs from the evaluation data used in Table 1.
Table 2 shows the impact of pretraining and fine-tuning (FT). PinRec-OC (the top row) is the baseline where we test the pretrained model as-is, without no fine-tuning. We show the relative lift of other methods, $\frac{\text{method}-PinRec\text{-}OC}{PinRec\text{-}OC}$. We evaluate different variants of fine-tuning:
  • {FT}: Fine-tuning with using surface-specific actions as additional positives (Eq. 4).
  • {FT, IN} (Impression as Negatives): Impressions only items are used as in-batch negatives during surface fine-tuning.
  • {FT, IP} (Impression as Positives): Impressions only items in the surface are treated similarly as engaged items.
We consistently see the strong gains from fine-tuning, supporting our claim that each surface benefits from tuning the model to learn its distinct user behavior patterns and user/item demographics. Performance improves further when we add impression-only items as in-batch negatives, since they act as hard “seen-but-not-chosen” counterexamples. In contrast, treating impression-only items as positives (IP) rewards exposure rather than intent, introducing noisy positives.
5.2.3. Outcome-Conditioning’s Impact on Fine-tuning
We further show that fine-tuning without outcome conditioning remains inferior to fine-tuning with outcome conditioning, supporting our claim that outcome conditioning is essential for fully optimizing GR performance on each surface. The last row of Table 2 reports the relative change of the fine-tuned unconditioned model PinRec-UC compared to the pretrained non-fine-tuned outcome-conditioned model PinRec-OC: $\frac{PinRec\text{-}UC\{FT,IN\}-PinRec\text{-}OC}{PinRec\text{-}OC}$. For all surfaces, PinRec-UC significantly underperforms PinRec-OC even after fine-tuning, indicating that fine-tuning alone cannot compensate for the absence of outcome conditioning. Together, these results show that outcome conditioning is necessary, and that combining it with fine-tuning yields the best performance.
Table 2. Comparison of fine-tuning PinRec-OC variants (recall @ 10) across major surfaces at Pinterest.
Model Home Feed Related Pins Search
PinRec-OC (no FT) 0.0% 0.0% 0.0%
PinRec-OC {FT} +4.1% +2.1% +2.5%
PinRec-OC {FT, IN} +4.5% +2.0% +2.5%
PinRec-OC {FT, IP} +4.1% -1.0% -4.3%
PinRec-UC {FT, IN} -8.1% -5.5% -9.4%
Table 3. Percentage lift in recall@10 for PinRec-OC over PinRec-UC when matching outcomes and desired actions.
Surface Outcome Recall Lift
Home Feed Save +1.9%
Outbound Click +6.2%
Search Product Outbound Click +16%
All other actions +8%
5.2.4. Outcome-Conditioning’s Impact on Metrics
@@
We assess whether conditioning on a desired action (e.g., “save”) leads to outputs that elicit that action. If our recommendations successfully control outcomes, we expect higher offline recall for the conditioned action.
With PinRec-OC{FT, IN} models, we allocate the entire budget to a single action and run recall only for the action (Table 3). We see larger improvements for rarer outcomes. For example, product outbound click is a rare but important search action, since Pinterest wants users to click through to products they intend to shop. In this case, PinRec-OC delivers a 16% lift.
@@ 6. Online A/B Experiments
We perform online A/B experiments to evaluate PinRec as a Candidate Generator (CG), measuring its impact on user engagement and its ability to tune the trade-off between business metrics. The items recommended by PinRec are passed through (unmodified) ranking and blending stages on each surface, alongside other Candidate Generators (CGs), to produce the final results shown to users. We tested on the three major surfaces: Home Feed, Search and Related Pins.
6.1. Trading Off Objectives: A Pareto Analysis
To validate that PinRec can efficiently control trade-offs between business objectives, we run five PinRec-OC configurations with varying action budgets $\mathcal{B}$. We allocate a budget of $\mathcal{B}$ to outcome A and 11-$\mathcal{B}$ to the other outcome. Figure 5 shows the resulting Pareto frontier for the two outcome rates (normalized by impressions) produced by the PinRec candidate generator alone; absolute values are omitted due to non-public information. These results show that PinRec enables the business to balance multiple metrics by tuning the trade-off between objectives. Additional results are shown in Figure 6.
(a) PinRec-OC Search Goals
(b) PinRec-OC HF Goals
Figure 5. Visualization of Pareto fronts for online engagement rates by outcome in surface-specific PinRec variants.
6.2. Surface A/B Tests